OpenStudy (96mertcank96):

Multiply -3-2i by is conjugate

OpenStudy (anonymous):

\[(a+bi)(a-bi)=a^2+b^2\] always in your case \(a=-3,b=-2\)

OpenStudy (anonymous):

ignore the minus signs ignore the \(i\) compute \(3^2+2^2\) and you are done

OpenStudy (triciaal):

to find the conjugate "change the sign before the term with i" i^2 = -1